OpenStudy (hba):

Center of the circle represented by the equation 5x^2+5y^2+24x+36y+10=0

OpenStudy (anonymous):

Use Completion Of Square Method..

OpenStudy (hba):

What ?

OpenStudy (anonymous):

Nothing.. I said use your method to solve it..

OpenStudy (hba):

Which method dude ?

OpenStudy (hba):

Hints please @waterineyes

OpenStudy (anonymous):

There you have open the square, here you will close the square.. I mean, write that in the form of square.. \((....)^2\)

OpenStudy (hba):

@waterineyes ??

OpenStudy (hba):

Let me try :(

OpenStudy (anonymous):

Can I give you some other example??

OpenStudy (anonymous):

Firstly make the coefficient of \(x^2\) and \(y^2\) as 1..

OpenStudy (anonymous):

Try first..

OpenStudy (hba):

I am comparing it let me do it dont answer

OpenStudy (hba):

Or i will keeeel you.

OpenStudy (hba):

@waterineyes (-g,-f)=(-12/5,-18/5)

OpenStudy (anonymous):

Wait, let me check..

OpenStudy (anonymous):

And you are right..

OpenStudy (hba):

Thanks A lot :)

OpenStudy (hba):

What will be the radius of this equation ? @waterineyes

OpenStudy (hba):

@hartnn Is it r=root(g^2+f^2-c)=0

OpenStudy (anonymous):

Now, solve for all the constants there..

OpenStudy (hba):

How am i going to survive without a calculator ?

OpenStudy (anonymous):

\[(x+\frac{12}{5})^2 + (y+\frac{18}{5})^2 - (\frac{12}{5})^2 - (\frac{18}{5})^2 +10 = 0\]

OpenStudy (anonymous):

Just solve the constant terms here..

hartnn (hartnn):

r=root(g^2+f^2-c) is the direct formula....if u need to show work, follow water.

OpenStudy (hba):

I do not need to show work lol,i just need to get the answer right :D

OpenStudy (anonymous):

What is c here?? 10 @hartnn

OpenStudy (anonymous):

Simple, then follow @hartnn Ha ha ha..

OpenStudy (hba):

I am using that formula lol :D

OpenStudy (hba):

And c = 10/5

OpenStudy (anonymous):

And I am not typing.. Ha ha ha ha...

OpenStudy (hba):

@waterineyes Common sense is not so common :)

OpenStudy (anonymous):

See, @hartnn has used there c = 10. Ha ha ha..

hartnn (hartnn):

c=10 here.

OpenStudy (anonymous):

So, we don't divide it here by 5 to c??

OpenStudy (hba):

Haha @hartnn you are wrong :(

hartnn (hartnn):

ok, my bad....i thot of different general equation....

OpenStudy (hba):

Anyways the answer is root(418)/5

OpenStudy (hba):

@hartnn ok fine.

OpenStudy (anonymous):

hartnn is taking time to think if it was wrong or right..

OpenStudy (hba):

@waterineyes You are confusing people today lol

OpenStudy (anonymous):

So, c = 2.. @hba now you cannot refuse here.. Ha ha ha..

hartnn (hartnn):

lets move on guys :P

OpenStudy (anonymous):

I am not confusing anybody, but confirming..

OpenStudy (hba):

Yeah lol

OpenStudy (hba):

Now as hartnn said move on :)

OpenStudy (anonymous):

Otherwise, I will explode your fuse.,. ha ha ha..

OpenStudy (anonymous):

How can I?? I am sitting on my chair.. Ha ha ha..

OpenStudy (hba):

From the question lol :D

OpenStudy (hba):

I thought i deserved a medal here :(

OpenStudy (anonymous):

Okay, this looks fine..

OpenStudy (anonymous):

@hartnn give your medal to @hba

OpenStudy (anonymous):

Okay, I give you now..

OpenStudy (anonymous):

I wish I could have more medals.. So that I can give them to me only.. Ha ha ha ha..